# Equivariant Machine Learning on Graphs with Nonlinear Spectral Filters

### v1

@@ 5.2 Node Classification on Filtered Chameleon and Squirrel Datasets in Dense Split Setting
Recently, the work in [Platonov et al., 2023] identified the presence of many duplicate nodes across the train, validation, and test splits in the dense split setting of the Chameleon and Squirrel [Pei et al., 2020]. This results in train-test data leakage, causing GNNs to inadvertently fit the test splits during training, thereby making performance results on Chameleon and Squirrel less reliable. To further validate the performance of our Node-level NLSFs on these datasets in the dense split setting, we use both the original and filtered versions of Chameleon and Squirrel, which do not contain duplicate nodes, as suggested in [Platonov et al., 2023]. We use the same random splits as in [Platonov et al., 2023], dividing the datasets into 48% for training, 32% for validation, and 20% for testing.
Tab. 2 presents the classification performance comparison between the original and filtered Chameleon and Squirrel. The baseline results are taken from [Platonov et al., 2023], and we include the following competitive models: ResNet+SGC [Platonov et al., 2023], ResNet+adj [Platonov et al., 2023], GCN [Kipf and Welling, 2016], GPRGNN [Chien et al., 2020], FSGNN [Maurya et al., 2022], GloGNN [Li et al., 2022], and FAGCN [Bo et al., 2021]. The detailed comparisons are in App. C. We see in the table that the $\mathtt{att}$-Node-level NLSFs consistently outperform the competing baselines on both the original and filtered datasets. $\mathtt{att}$-Node-level NLSFs demonstrate less sensitivity to node duplicates and exhibit stronger generalization ability, further validating the reliability of the Chameleon and Squirrel datasets in the dense split setting. We note that compared to the dense split setting, the sparse split setting in Tab. LABEL:tab:node_classification_accuracy is more challenging, resulting in lower classification performance. A similar trend of significant performance difference between the two settings of Chameleon and Squirrel is also observed in [He et al., 2022].
@@ 5.3 Graph Classification
We further illustrate the power of NLSFs on eight graph classification benchmarks [Kersting et al., 2016]. Specifically, we consider five bioinformatics datasets [Borgwardt et al., 2005, Kriege and Mutzel, 2012, Shervashidze et al., 2011]: MUTAG, PTC, NCI1, ENZYMES, and PROTEINS, where MUTAG, PTC, and NCI1 are characterized by discrete node features, while ENZYMES and PROTEINS have continuous node features. Additionally, we examine three social network datasets: IMDB-B, IMDB-M, and COLLAB. The unattributed graphs are augmented by adding node degree features following [Errica et al., 2019]. For more details of these datasets, see App. B.
In graph classification tasks, a readout function is used to summarize node representations for each graph. The final graph-level representation is obtained by aggregating these node-level summaries and is then fed into an MLP with a (log)softmax layer to perform the graph classification task. We compare our NLSFs with two kernel-based approaches: GK [Shervashidze et al., 2009] and WL [Shervashidze et al., 2011], as well as nine GNNs: GCN [Kipf and Welling, 2016], GAT [Veličković et al., 2017], SAGE [Hamilton et al., 2017], ChebNet [Defferrard et al., 2016a], ChebNetII [He et al., 2022], CayleyNet [Levie et al., 2018], APPNP [Klicpera et al., 2019], GPRGNN [Chien et al., 2020], and ARMA [Bianchi et al., 2021]. Additionally, we consider the hierarchical graph pooling model DiffPool [Ying et al., 2018]. For dataset splitting, we apply the random split following [Velickovic et al., 2019, Ying et al., 2018, Ma et al., 2019, Zhu et al., 2021b], using 80% for training, 10% for validation, and 10% for testing. This random splitting process is repeated 10 times, and we report the average performance along with the standard deviation. For the baseline algorithms, we use the source code released by the authors and optimize their hyperparameters using Optuna [Akiba et al., 2019]. We fine-tune the hyperparameters of each model to achieve the highest possible accuracy. Detailed hyperparameter settings for both the baselines and our method are provided in App. B.
@@ 6 Summary
Limitation and Future Work. One limitation of NLSFs is that, when deployed on large graphs, they only depend on the leading eigenvectors of the GSO. In this paper, we considered the low-frequency eigenvectors of graph Laplacians. However, important information can also lie within the high frequencies or in any other band. In future work, we plan to explore NLSFs that are sensitive to eigenvalues that lie within a set of bands of interest, which can be adaptive to the graph.
@@ Acknowledgments
Appendix A Theoretical Analysis

### v2

@@ 5.2 Node Classification on Filtered Chameleon and Squirrel Datasets in Dense Split Setting
$\mathtt{att}$-Node-level NLSFs 79.84$\pm$1.2 42.53$\pm$1.5 68.17$\pm$1.9 42.66$\pm$1.7
Recently, the work in [77] identified the presence of many duplicate nodes across the train, validation, and test splits in the dense split setting of the Chameleon and Squirrel [75]. This results in train-test data leakage, causing GNNs to inadvertently fit the test splits during training, thereby making performance results on Chameleon and Squirrel less reliable. To further validate the performance of our Node-level NLSFs on these datasets in the dense split setting, we use both the original and filtered versions of Chameleon and Squirrel, which do not contain duplicate nodes, as suggested in [77]. We use the same random splits as in [77], dividing the datasets into 48% for training, 32% for validation, and 20% for testing.
Tab. 2 presents the classification performance comparison between the original and filtered Chameleon and Squirrel. The baseline results are taken from [77], and we include the following competitive models: ResNet+SGC [77], ResNet+adj [77], GCN [49], GPRGNN [16], FSGNN [69], GloGNN [62], and FAGCN [8]. The detailed comparisons are in App. F. We see in the table that the $\mathtt{att}$-Node-level NLSFs consistently outperform the competing baselines on both the original and filtered datasets. $\mathtt{att}$-Node-level NLSFs demonstrate less sensitivity to node duplicates and exhibit stronger generalization ability, further validating the reliability of the Chameleon and Squirrel datasets in the dense split setting. We note that compared to the dense split setting, the sparse split setting in Tab. LABEL:tab:node_classification_accuracy is more challenging, resulting in lower classification performance. A similar trend of significant performance difference between the two settings of Chameleon and Squirrel is also observed in [41].
@@ 5.3 Graph Classification
We further illustrate the power of NLSFs on eight graph classification benchmarks [47]. Specifically, we consider five bioinformatics datasets [10, 53, 86]: MUTAG, PTC, NCI1, ENZYMES, and PROTEINS, where MUTAG, PTC, and NCI1 are characterized by discrete node features, while ENZYMES and PROTEINS have continuous node features. Additionally, we examine three social network datasets: IMDB-B, IMDB-M, and COLLAB. The unattributed graphs are augmented by adding node degree features following [26]. For more details of these datasets, see App. E.
In graph classification tasks, a readout function is used to summarize node representations for each graph. The final graph-level representation is obtained by aggregating these node-level summaries and is then fed into an MLP with a (log)softmax layer to perform the graph classification task. We compare our NLSFs with two kernel-based approaches: GK [87] and WL [86], as well as nine GNNs: GCN [49], GAT [93], SAGE [39], ChebNet [22], ChebNetII [41], CayleyNet [61], APPNP [50], GPRGNN [16], and ARMA [5]. Additionally, we consider the hierarchical graph pooling model DiffPool [101]. For dataset splitting, we apply the random split following [94, 101, 67, 104], using 80% for training, 10% for validation, and 10% for testing. This random splitting process is repeated 10 times, and we report the average performance along with the standard deviation. For the baseline algorithms, we use the source code released by the authors and optimize their hyperparameters using Optuna [1]. We fine-tune the hyperparameters of each model to achieve the highest possible accuracy. Detailed hyperparameter settings for both the baselines and our method are provided in App. E.
@@ 6 Summary
Limitation and Future Work. One limitation of NLSFs is that, when deployed on large graphs, they only depend on the leading eigenvectors of the GSO and their orthogonal complement. However, important information can also lie within any other band. In future work, we plan to explore NLSFs that are sensitive to eigenvalues that lie within a set of bands of interest, which can be adaptive to the graph.
@@ Acknowledgments
Appendix A Non-linear Activation Functions Break the Functional Symmetry
We offer a construction for complex-valued signals and first order derivative as the graph Laplacian, which is easy to extend to the real case and second order derivative. Consider a circular 1D grid with $100$ nodes and the first order central difference as the Laplacian. Here, the Laplacian eigenvectors are the standard Fourier modes. In this case, one can see that a graph functional shift is any operator that is diagonal in the frequency domain and multiplies each frequency by any complex number with the unit norm. Consider the nonlinearity that takes the real part of the signal and then applies ReLU, which we denote in short by ReLU. Consider a graph signal $x=(e^{i\pi 10n/100}+e^{i\pi 20n/100})_{n=0}^{99}$. We consider a graph functional shift $S$ that shifts frequencies 10 and 20 at a distance of 5, and every other frequency is not shifted. Namely, frequency 10 is multiplied by $e^{-i\pi 50/100}=-i$, frequency 20 by $e^{-i\pi=-i\pi 100/100}=-1$, and every other frequency is multiplied by 1. Consider also the classical shift $D$ that translates the whole signal by $5$ uniformly. Since $x$ consists only of the frequencies 10 and 20, it is easy to see that $Sx=Dx$. Hence, $\text{ReLU}(Sx)=\text{ReLU}(Dx)=D(\text{ReLU}(x))$. Conversely, if we apply $\text{ReLU}(x)$ and only then shift, note that $\text{ReLU}(x)$ consists of many frequencies in addition to 10 and 20. For example, by nonnegativity of ReLU, $\text{ReLU}(x)$ has a nonzero DC component (zeroth frequency). Now, $S(\text{ReLU}(x))$ only translates the 10 and 20 frequencies, so we have $S(\text{ReLU}(x))\neq D(\text{ReLU}(x))=\text{ReLU}(S(x))$.
Appendix B Illustrating Functional Translations
Here we present an additional discussion and illustrations on the new notions of symmetry.
B.1 Functional Translation of a Gaussian Signal with Different Speeds at Different Frequencies
To illustrate the concepts of relaxed symmetry and functional translation, we present a toy example involving the classical translation and a functional translation of a Gaussian signal on a 2D grid with a standard deviation of 1.
The classical translation involves moving the entire signal uniformly across the grid. This uniform movement can be represented as $I\textquoteright(x,y)=I(x-t_{x},y-t_{y})$, where $t_{x}$ and $t_{y}$ are the translation amounts in the $x$ and $y$ directions, respectively, and $x-t_{x}$ and $y-t_{y}$ are circular translations, namely, subtractions modulo the size of the circular grid. For simplicity, we consider $t_{x}=t_{y}=t$. For instance, if $t=5$, the Gaussian is shifted by 5 units in both the $x$ and $y$ directions. Fig. 3 top-row shows the classical translation for $t=0$, $t=5$, $t=10$, and $t=15$. We see that every part of the signal shifts at the same rate and direction, preserving the overall shape of the Gaussian signal. Note that classical translation is equivalent to modulation of the frequency domain, i.e. $\widehat{I}\textquoteright(u,v)=\widehat{I}(u,v)e^{-i2\pi(ut_{x}+vt_{y})}$, where $\widehat{I}(u,v)$ is the Fourier transform of $I(x,y)$. Therefore, we can view the classical translation as a specific type of functional translation.
Next, we illustrate a functional translation that is based on a frequency-dependent movement. Specifically, the Gaussian signal is decomposed into low and high-frequency components based on two indicator band-pass filters. In our example, the translation parameter $t$ is different for the low and high-frequency components: low frequencies are shifted by $t_{low}$ while high frequencies are shifted by $t_{high}$. This functional translation is defined via modulations in the frequency domain given by $\widehat{I}_{low}^{\prime}(u,v)=\widehat{I}_{low}(u,v)e^{-i2\pi(ut_{x,low}+vt_{y,low})}$ for low-frequency components and $\widehat{I}_{high}^{\prime}(u,v)=\widehat{I}_{high}(u,v)e^{-i2\pi(ut_{x,high}+vt_{y,high})}$ for high-frequency components. The combined functionally translated signal in the frequency domain is then $\widehat{I}^{\prime}=(\widehat{I}_{low}^{\prime},\widehat{I}_{high}^{\prime})$. Fig. 3 bottom-row demonstrates the functional translation for $t=0$, $t=5$, $t=10$, and $t=15$. We observe that the low-frequency components (smooth parts) of the signal move at one speed, while high-frequency components move at another. This demonstrates that functional symmetries are typically more rich than domain symmetries.
Figure 3: Comparison of classical and functional translation of a Gaussian signal. Top Row (Classical Translation): The Gaussian signal moves uniformly across the grid without changing shape. Bottom Row (Functional Translation): The Gaussian signal translates as low-frequency components move at different speeds than high-frequency components, demonstrating relaxed symmetry.
B.2 Robust Graph Functional Shifts in Image Translation and MNIST Classification on Perturbed Grids
We present another toy example to illustrate that functional translations are more stable than hard symmetries of the graph, namely, graph automorphisms (isomorphisms from the graph to itself). Automorphisms are another analog to translations on general graphs, which competes with our notion of functional shifts. Consider as an example the standard 2D circular grid (discretizing the torus). The automorphisms of the grid include all translations. However, this notion of hard symmetry is very sensitive to graph perturbations. If one adds or deletes even one edge in the graph, the automorphism group becomes much smaller and does not contain the translations anymore.
In contrast, we claim that functional shifts are not so sensitive to graph perturbations. To demonstrate this empirically, we conducted the following toy experiment. We add a Gaussian noise to the edge weights of the 2D grid to create a perturbed graph. Given a standard domain shift, we optimize the coefficients of a functional shift so it is as close as possible to the classical domain shift of the clean grid in Frobenius error. Fig. 4 presents an example of classically and functionally shifted image of the digit 5. The original digit (left), a classical domain translation of the clean grid (middle), and a functional translation constructed from the perturbed graph (right) to match the classical translation. We see that standard translations can be approximated by a graph functional shift on a perturbed grid.
We further demonstrate the robustness to graph perturbations of NLSFs trained on MNIST classification. The experimental setup follows previous work in [22, 71, 61]. We compare the NLSF on the clean grid (i.e., without Gaussian noise) to the NLSF on the perturbed grid. Tab. 4 presents the classification accuracy of the NLSF. We see that the classification performance is almost unaffected by the perturbation, indicating that NLSFs on the perturbed grid can roughly express CNN-like operations (translation equivariant functions).
Figure 4: Approximate a standard translation by functional translation on a perturbed graph.
MNIST Perturbed MNIST
Ours 99.19 99.16
Table 4: Classification accuracy on the MNIST and perturbed MNIST using NLSFs.
Appendix C Special Cases of NLSFs
We present two special cases of NLSFs as follows.
C.1 Diagonal NLSFs
In Sec. 3, we introduced the NLSFs with the output dimension $\widetilde{d}$, which is a tunable hyperparameter. Here, we present a special case when $\widetilde{d}=d$ such that the multiplication and division in synthesis are operated diagonally. Specifically, the diagonal analysis and synthesis in the spectral index case and in the filter bank case are defined respectively by
$\mathcal{A}^{\text{ind, diag}}({\bm{\Delta}},{\mathbf{X}})$ $=\Big{(}\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}\Big{)}_{j=1}^{J+1}\in{\mathbb{R}}^{(J+1)\times d}\text{ \ and \ }$ (7)
$\mathcal{A}^{\text{val, diag}}({\bm{\Delta}},{\mathbf{X}})$ $=\Big{(}\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\Big{)}_{j=1}^{K+1}\in{\mathbb{R}}^{(K+1)\times d}$ (8)
and
$\mathcal{S}^{\text{ind, diag}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})$ $=\sum_{j=1}^{J+1}\mathbf{r}_{j}\odot\ \frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\text{\ and \ }$ (9)
$\mathcal{S}^{\text{val, diag}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})$ $=\sum_{j=1}^{K+1}\mathbf{r}_{j}\odot\frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e},$ (10)
where the product and division are element-wise along the channel direction. That is, $\mathbf{r}_{j}\odot\ \frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}=\left[{r_{j}}_{1}\frac{\mathbf{P}_{j}{\mathbf{X}}_{:,1}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}_{:,1}\right\rVert_{2}^{a}+e},\ldots,{r_{j}}_{d}\frac{\mathbf{P}_{j}{\mathbf{X}}_{:,d}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}_{:,d}\right\rVert_{2}^{a}+e}\right]$ in the spectral index case (resp. $\mathbf{r}_{j}\odot\frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}=\left[{r_{j}}_{1}\frac{g_{j}({\bm{\Delta}}){\mathbf{X}}_{:,1}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}_{:,1}\right\rVert_{2}^{a}+e},\ldots,{r_{j}}_{d}\frac{g_{j}({\bm{\Delta}}){\mathbf{X}}_{:,d}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}_{:,d}\right\rVert_{2}^{a}+e}\right]$ in the filter bank case). Here, $\mathbf{R}=(\mathbf{r}_{j})_{j=1}^{J+1}\in{\mathbb{R}}^{(J+1)\times d}$ in the spectral index case (resp. $\mathbf{R}=(\mathbf{r}_{j})_{j=1}^{K+1}\in{\mathbb{R}}^{(K+1)\times d}$ in the filter bank case) are the spectral coefficients to be synthesized and $a,e$ are as before.
For Node-level diag-NLSFs, we define the Index diag-NLSFs and Value diag-NLSFs as follows
$\Theta_{\text{ind, diag}}({\bm{\Delta}},{\mathbf{X}})$ $=\mathcal{S}^{(\text{ind, diag})}_{{\bm{\Delta}},{\mathbf{X}}}(\Psi(\mathcal{A}^{(\text{ind})}_{{\bm{\Delta}}}({\mathbf{X}})))=\sum_{j=1}^{J+1}\left[\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J+1}\right]_{j}\ \frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e},$ (11)
$\Theta_{\text{val, diag}}({\bm{\Delta}},{\mathbf{X}})$ $=\mathcal{S}^{(\text{val, diag})}_{{\bm{\Delta}},{\mathbf{X}}}(\Psi(\mathcal{A}^{(\text{val})}_{{\bm{\Delta}}}({\mathbf{X}})))=\sum_{j=1}^{K+1}\left[\Psi\left(\left\lVert g_{i}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{K+1}\right]_{j}\ \frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e},$ (12)
where $\Psi:{\mathbb{R}}^{(J+1)d}\rightarrow{\mathbb{R}}^{(J+1)d}$ in $\Theta_{\text{ind, diag}}$ and $\Psi:{\mathbb{R}}^{(K+1)d}\rightarrow{\mathbb{R}}^{(K+1)d}$ in $\Theta_{\text{val, diag}}$. To adjust the feature output dimension at each node, we apply an MLP with shared weights to all nodes after the NLSF. We present the empirical study of diag-NLSFs in App. F.7.
C.2 Leading NLSFs
In Sec. 3.3, we introduce the NLSFs with the orthogonal complement. Specifically, the $(J+1)$-th filter in the Index NLSFs is given by $\mathbf{P}_{J+1}=\mathbf{I}-\sum_{j=1}^{J}\mathbf{P}_{j}$, and the $(K+1)$-th filter in Vale NLSFs is defined as $g_{K+1}(\mathbf{\Delta})=\mathbf{I}-\sum_{j=1}^{K}g_{j}(\mathbf{\Delta})$. To explore the effects of the orthogonal complement, we focus on the leading NLSFs in both the Index NLSFs and Vale NLSFs, considering only the first $J$ and $K$ filters without including their orthogonal complements. In this case, the analysis and synthesis in the spectral index case and in the filter bank case are respectively given by
$\begin{split}&\mathcal{A}^{\text{ind, lead}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})=\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\in{\mathbb{R}}^{Jd}\text{\ and \ }\\ &\mathcal{A}^{\text{val, lead}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})=\left(\left\lVert g_{i}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{K}\in{\mathbb{R}}^{Kd},\end{split}$ (13)
and
$\begin{split}&\mathcal{S}^{\text{ind, lead}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})=\left[\frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\right]_{j=1}^{J}\mathbf{R}\text{\ and \ }\\ &\mathcal{S}^{\text{val, lead}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})=\left[\frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\right]_{j=1}^{K}\mathbf{R},\end{split}$ (14)
where $\mathbf{P}_{j}$ and $g_{j}({\bm{\Delta}})$ are defined as in Sec. 3.2. The lead-NLSFs are then defined by
$\begin{split}\Theta_{\text{ind, lead}}({\bm{\Delta}},{\mathbf{X}})&=\mathcal{S}^{(\text{ind, lead})}_{{\bm{\Delta}},{\mathbf{X}}}(\Psi_{\text{ind}}(\mathcal{A}^{(\text{ind})}_{{\bm{\Delta}}}({\mathbf{X}})))\\ &=\left[\frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\right]_{j=1}^{J}\left[\Psi_{\text{ind}}\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)\right]_{i=1}^{J},\end{split}$ (15)
$\begin{split}\Theta_{\text{val, ind}}({\bm{\Delta}},{\mathbf{X}})&=\mathcal{S}^{(\text{val, ind})}_{{\bm{\Delta}},{\mathbf{X}}}(\Psi_{\text{val}}(\mathcal{A}^{(\text{val})}_{{\bm{\Delta}}}({\mathbf{X}})))\\ &=\left[\frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\right]_{j=1}^{K}\left[\Psi_{\text{val}}\left(\left\lVert g_{i}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\right)\right]_{i=1}^{K},\end{split}$ (16)
where $\Psi_{\text{ind}}:{\mathbb{R}}^{Jd}\rightarrow{\mathbb{R}}^{Jd\times J\widetilde{d}}$, $\Psi_{\text{val}}:{\mathbb{R}}^{Kd}\rightarrow{\mathbb{R}}^{Kd\times K\widetilde{d}}$, and $\widetilde{d}$ is the output dimension. To adjust the feature output dimension, we apply an MLP with shared weights to all nodes after the NLSF.
Similar to App. C.1, when considering $\widetilde{d}=d$ such that the multiplication and division in synthesis are operated diagonally, we have the lead-diag-NLSFs defined by
$\begin{split}\Theta_{\text{ind, lead, diag}}({\bm{\Delta}},{\mathbf{X}})&=\mathcal{S}^{(\text{ind, lead, diag})}_{{\bm{\Delta}},{\mathbf{X}}}(\Psi(\mathcal{A}^{(\text{ind})}_{{\bm{\Delta}}}({\mathbf{X}})))\\ &=\sum_{j=1}^{J}\left[\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\right]_{j}\ \frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e},\end{split}$ (17)
$\begin{split}\Theta_{\text{val, lead, diag}}({\bm{\Delta}},{\mathbf{X}})&=\mathcal{S}^{(\text{val, lead, diag})}_{{\bm{\Delta}},{\mathbf{X}}}(\Psi(\mathcal{A}^{(\text{val})}_{{\bm{\Delta}}}({\mathbf{X}})))\\ &=\sum_{j=1}^{K}\left[\Psi\left(\left\lVert g_{i}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{K}\right]_{j}\ \frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e},\end{split}$ (18)
where $\Psi:{\mathbb{R}}^{Jd}\rightarrow{\mathbb{R}}^{Jd}$ in $\Theta_{\text{ind, lead, diag}}$ and $\Psi:{\mathbb{R}}^{Kd}\rightarrow{\mathbb{R}}^{Kd}$ in $\Theta_{\text{val, lead, diag}}$. We present the empirical study of lead-NLSFs and lead-diag-NLSFs in App. F.7.
Appendix D Theoretical Analysis